# Robust Learning of Multi-index Models via Iterative Subspace Approximation

### v1

@@ 2.2 Finding a Relevant Direction
If $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{{\mathcal{S}}}(\mathbf{x})\neq y]>(C+1)\mathrm{OPT}+5\epsilon$, then there exists a subset $\mathcal{T}\subseteq\mathcal{S}$ such that $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[S]\geq\epsilon$; and for all $i\in[K],S\in\mathcal{T}$, it holds that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq i\mid\mathbf{x}\in S]>C\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq y\mid\mathbf{x}\in S]+\epsilon$.
@@
Define $\mathcal{T}$ to be all sets $S\subseteq\mathcal{S}$ such that $\operatorname*{\mathbf{Pr}}[i\neq y\mid\mathbf{x}\in S]>(C+1)\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq y\mid\mathbf{x}\in S]+2\epsilon$, for all $i\in[K],S\in\mathcal{T}$. Assume for the sake of contradiction that $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[S]\leq 2\epsilon$. We have that
@@
$\leq\sum_{S\not\in\mathcal{T}}\min_{\widehat{y}\in[K]}\left((C+1)\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y\mid\mathbf{x}\in S]+2\epsilon\right)\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[S]+3\epsilon$
@@
where in the first equality, we used the fact that at most an $\epsilon$ fraction of the mass is outside the discretization (see A.1). This leads to a contradiction and hence $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[S]>2\epsilon$.
Furthermore, note that by Lemma 2.10 there exists $\mathcal{T}^{\prime}\subseteq\mathcal{T}$ such that $\sum_{S\in\mathcal{T}^{\prime}}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[S]>\epsilon$ and for all $i\in[K],S\in\mathcal{T}^{\prime}$, it holds that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[i\neq y\mid\mathbf{x}\in S]>(C+1)\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f_{\mathcal{G}_{S}}(x)\neq y\mid\mathbf{x}\in S]+\epsilon$. Since $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[i\neq y\mid\mathbf{x}\in S]\leq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq i\mid\mathbf{x}\in S]+\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq y\mid\mathbf{x}\in S]$ we have that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq i\mid\mathbf{x}\in S]>C\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq y\mid\mathbf{x}\in S]+\epsilon$ for all $S\in\mathcal{T}^{\prime}$, which completes the proof. ∎
@@
Let $S\in\mathcal{S}$ be a subset of the partition such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq i\mid\mathbf{x}\in S]>C\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq y\mid\mathbf{x}\in S]+\epsilon$. Then there exists some $i\in[K]$ and a unit vector $\mathbf{w}\in W$ for which $(\widehat{\mathbf{w}}^{(i,j)})^{\perp V}\cdot\mathbf{u}^{(S,i)}\geq\mathrm{poly}(1/K,\epsilon)$.
@@
First, notice that the output label of $f_{\mathcal{G}_{S}}$ is only determined by the projection at the orthogonal complement of $V$, i.e.,
@@
Furthermore, note that the condition whether $\mathbf{x}$ belongs to $S$ depends only on the projection of $\mathbf{x}^{V}$, hence without loss of generality it can written as $\mathds{1}(\mathbf{x}^{V}\in S)$. Thus, we have that
@@
$=\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq y]\;,$
where $D_{V^{\perp}}$ is defined to be the marginal obtained by averaging $D$ over $V$ conditioned on $S$. Similarly, we have that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq i]>C\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq y]+\epsilon$ for all $i\in[K]$. Moreover, by the properties of the standard Gaussian distribution, we have that the $\mathbf{x}$-marginal of $D_{V^{\perp}}$ is a standard Gaussian.
As a result, by applying Proposition 2.4 for $f_{\mathcal{G}_{S}}$, we have that there exist $i,j\in[K],i\neq j$ such that $(\widehat{\mathbf{w}}^{(i,j)})^{\perp V}\cdot\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[\mathbf{x}\mathds{1}[y=i]]\geq\mathrm{poly}(1/K,\epsilon)$. Furthermore, note that by averaging over $V$ conditioning on $S$ we have that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{x}^{\perp V}\mathds{1}[y=i]\mid\mathbf{x}\in S]=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[\mathbf{x}\mathds{1}[y=i]]$. Therefore, there exist $i,j\in[K],i\neq j$, such that $(\widehat{\mathbf{w}}^{(i,j)})^{\perp V}\cdot\mathbf{u}^{(S,l)}\geq\mathrm{poly}(1/K,\epsilon)$ for some $l\in[K]$.
@@
Let $p(\epsilon,1/K)$ be a sufficiently small polynomial in $\epsilon$ and $1/K$. If $N=d(k/\epsilon^{\prime})^{Ck}(K/\epsilon)^{C}\log(1/\delta)$ for a sufficiently large universal constant $C>0$, then with probability $1-\delta$ we have that $\left\|\widehat{\mathbf{u}}^{(S,i)}-{\mathbf{u}}^{(S,i)}\right\|\leq{p(\epsilon,1/K)}$ and $\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]\geq\operatorname*{\mathbf{Pr}}_{D}[S]/2$ for all $S\in\mathcal{S},i\in[K]$.
@@
First note that by the definition of an approximating partition (see Definition 2.2) and A.1, for all $S\in\mathcal{S}$ we have that $\operatorname*{\mathbf{Pr}}_{D}[S]{=(\epsilon^{\prime}/k)^{\Omega(k)}}$ and $\lvert\mathcal{S}\rvert{=(k/\epsilon^{\prime})^{O(k)}}$. Therefore, by union bound and Hoeffding’s inequality, it holds that if $N\geq{(k/\epsilon^{\prime})^{Ck}}\log(\lvert\mathcal{S}\rvert/\delta)={(k/\epsilon^{\prime})^{O(k)}}\log(1/\delta)$ for a sufficiently large constant $C>0$, then with probability at least $1-\delta$ we have that $\lvert\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]-\operatorname*{\mathbf{Pr}}_{D}[S]\rvert\leq\operatorname*{\mathbf{Pr}}_{D}[S]/2$ for all $S\in\mathcal{S}$. Hence, it is true that $\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]\geq\operatorname*{\mathbf{Pr}}_{D}[S]/2$, i.e., the number of samples that fall in each set $S\in\mathcal{S}$ is at least $N\operatorname*{\mathbf{Pr}}_{D}[S]/2{=N(\epsilon^{\prime}/k)^{\Omega(k)}}$.
@@
Furthermore, note that conditioning on any $S$, the random vector $\mathbf{x}\mathds{1}(y=i)$ with $(\mathbf{x},y)\sim D$ is subgaussian in any unit direction with subgaussian norm at most a constant. Because, we can decompose $\mathbf{v}\cdot\mathbf{x}=\mathbf{v}\cdot\mathbf{x}^{V}+\mathbf{v}\cdot\mathbf{x}^{\perp V}$ with $\mathbf{v}\cdot\mathbf{x}^{V},\mathbf{v}\cdot\mathbf{x}^{\perp V}$ being independent, where the zero mean centering $\mathbf{v}\cdot\mathbf{x}^{V}$ is bounded in any direction by $\epsilon^{\prime}\sqrt{k}\leq\epsilon$ since we have conditioned on $S$ and $\mathbf{v}\cdot\mathbf{x}^{\perp V}$ is a standard gaussian since $S$ depends only on $\mathbf{x}^{V}$. Therefore, by the use of a standard covering argument it follows that if the number of samples that falls in each discretization set $S$ is at least $d\mathrm{poly}(K,1/\epsilon)\log(K/\delta)$ for a sufficiently large polynomial, then $\left\|\widehat{\mathbf{u}}^{(S,i)}-{\mathbf{u}}^{(S,i)}\right\|\leq{p(\epsilon,1/K)}$ for all $i\in[K]$.
@@
Combining the above results with the allong with union bound gives us that there exists an $N=d(k/\epsilon^{\prime})^{O(k)}(K/\epsilon)^{O(1)}\log(1/\delta)$ that suffices.
@@
Applying Claims 2.11 and 2.12, we have that if $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{{\mathcal{S}}}(\mathbf{x})\neq y]>(C+1)\mathrm{OPT}+\epsilon$ then there exists a subset $\mathcal{T}\subseteq\mathcal{S}$ of the $\epsilon^{\prime}$-approximating partition with $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}_{D}[S]=\Omega(\epsilon)$ such that for all $S\in\mathcal{T}$ we have that $(\widehat{\mathbf{w}}^{(i,j)})^{\perp V}\cdot\mathbf{u}^{(S,l)}\geq\mathrm{poly}(1/K,\epsilon)$ for some $i,j,l\in[K],i\neq j$. Moreover, if $N=d(k/\epsilon^{\prime})^{C^{\prime}k}(K/\epsilon)^{C^{\prime}}\log(1/\delta)$ for a sufficiently large universal constant $C^{\prime}>0$, using the concentration result of Claim 2.13, we have that the empirical moments of each set satisfy the same guarantees, i.e., $(\widehat{\mathbf{w}}^{(i,j)})^{\perp V}\cdot\widehat{\mathbf{u}}^{(S,l)}\geq\mathrm{poly}(1/K,\epsilon)$, for some $i,j,l\in[K],i\neq j,$ for all $S\in\mathcal{T}$. Furthermore, note that by Claim 2.13 we also have that $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]=\Omega(\epsilon)$.
@@
Therefore, there exists a subset $\mathcal{T}^{\prime}\subseteq\mathcal{T}$ and $i,j\in[K]$ with $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]\geq\Omega(\epsilon/K^{2})$ such that for all $S\in\mathcal{T}$ we have that $(\widehat{\mathbf{w}}^{(i,j)})^{\perp V}\cdot\widehat{\mathbf{u}}^{(S,l)}\geq\mathrm{poly}(1/K,\epsilon)$ for some $l\in[K]$. Hence, for some $i\neq j\in[K]$ we have that
@@
Moreover, note that $(\mathbf{u}^{(S,i)})^{\perp V}=\operatorname*{\mathbf{E}}_{\mathbf{x}^{\perp V}}[\mathbf{x}^{\perp V}\operatorname*{\mathbf{E}}_{\mathbf{x}^{V}}[\mathds{1}(y=i)\mid\mathbf{x}\in S]]$ as $S$ depends only on $\mathbf{x}^{V}$ and
@@
where we used the Cauchy-Schwarz inequality and the dual characterization of the $\ell_{2}$ norm. Therefore, by the triangle inequality, it follows that
@@
Hence, by Claim B.1, we have that there exists a unit eigenvector $\mathbf{v}$ of $\widehat{\mathbf{U}}$ with eigenvalue at least ${\mathrm{poly}(\epsilon/K)}$ for a sufficiently small polynomial, such that $(\widehat{\mathbf{w}}^{(i,j)})^{\perp V}\cdot\mathbf{v}\geq\mathrm{poly}(1/K,\epsilon)$, for some $i\neq j\in[K]$. Moreover, the number of such eigenvectors is at most $2K/\mathrm{poly}(1/K,\epsilon)=\mathrm{poly}(K,1/\epsilon)$, which completes the proof of Proposition 2.8.
Before proving our main result of this section, Theorem 2.1, we need to show that if we have learned a direction of our target model to sufficiently high accuracy, then we can assume, without loss of generality, that the model is defined only on the remaining directions. Specifically, we establish the following:

### v2

@@ 2.2 Finding a Relevant Direction
If $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{\mathcal{S}}(\mathbf{x})\neq y]>(C+1)\mathrm{OPT}+5\epsilon$, then there exists a subcollection $\mathcal{T}\subseteq\mathcal{S}$ satisfying $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[S]\geq\epsilon$; and for every $S\in\mathcal{T}$ we have $\min_{i\in[K]}\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq i\mid\mathbf{x}\in S]>C\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq y\mid\mathbf{x}\in S]+\epsilon$.
@@
Define $\mathcal{T}$ to be the collection of $S\subseteq\mathcal{S}$ such that $\min_{i\in[K]}\operatorname*{\mathbf{Pr}}[i\neq y\mid\mathbf{x}\in S]>(C+1)\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq y\mid\mathbf{x}\in S]+2\epsilon$. Assume for the sake of contradiction that $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[S]\leq 2\epsilon$. We have that
@@
$\leq\sum_{S\not\in\mathcal{T}}\left((C+1)\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y\mid\mathbf{x}\in S]+2\epsilon\right)\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[S]+3\epsilon$
@@
where in the first line we used the fact (see A.1) that the mass of points outside the discretization is at most $\epsilon$, in the second line we applied our definition of $\mathcal{T}$ and in the last inequality we used that the total error of $f$ is $\mathrm{OPT}$. This contradicts our hypothesis that $\operatorname*{\mathbf{Pr}}[h_{\mathcal{S}}(\mathbf{x})\neq y]>(C+1)\mathrm{OPT}+5\epsilon$. Hence, it must be that $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[S]>2\epsilon$.
Next, recall that the cube width $\epsilon^{\prime}$ is chosen to be less than $c\epsilon^{2}/(K^{2}\sqrt{k})$ for a sufficiently small constant $c$. Taking $c<1/(C+1)$, we can apply Lemma 2.10 and get that there exists a collection $\mathcal{T}^{\prime}\subseteq\mathcal{S}$ with $\sum_{S\in\mathcal{T}^{\prime}}\operatorname*{\mathbf{Pr}}[S]\geq 1-\epsilon$, such that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f(\mathbf{x})\neq f_{\mathcal{G}_{S}}(\mathbf{x})\mid\mathbf{x}\in S]\leq\epsilon/(C+1)$. Note that for $\mathcal{T}^{\prime\prime}=\mathcal{T}^{\prime}\cap\mathcal{T}$, it must hold that $\sum_{S\in\mathcal{T}^{\prime\prime}}\operatorname*{\mathbf{Pr}}[s]\geq\epsilon$. For all $S\in\mathcal{T}^{\prime\prime}$, it holds that $\min_{i\in[K]}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[i\neq y\mid\mathbf{x}\in S]>(C+1)\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f_{\mathcal{G}_{S}}(x)\neq y\mid\mathbf{x}\in S]+\epsilon$. Since,
$\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\Bigl{[}i\neq y\mid\mathbf{x}\in S\Bigr{]}\leq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\Bigl{[}f_{\mathcal{G}_{S}}(\mathbf{x})\neq i\mid\mathbf{x}\in S\Bigr{]}+\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\Bigl{[}f_{\mathcal{G}_{S}}(\mathbf{x})\neq y\mid\mathbf{x}\in S\Bigr{]}\;,$
we get that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq i\mid\mathbf{x}\in S]>C\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq y\mid\mathbf{x}\in S]+\epsilon$ for all $S\in\mathcal{T}^{\prime\prime}$, which completes the proof. ∎
@@
Let $S\in\mathcal{S}$ such that $\min_{i\in[K]}\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq i\mid\mathbf{x}\in S]>C\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq y\mid\mathbf{x}\in S]+\epsilon$. Then there exists $i,j,l\in[K]$ with $i\neq j$ such that $(\widehat{\mathbf{w}}^{(i,j)})^{\perp V}\cdot\mathbf{u}^{(S,l)}\geq\mathrm{poly}(1/K,\epsilon)$.
@@
First, notice that the output label of $f_{\mathcal{G}_{S}}$ is only determined by the projection onto the orthogonal complement of $V$, i.e.,
@@
Furthermore, observe that whether $\mathbf{x}$ belongs to $S$ depends only on its projection onto $V$, hence without loss of generality it can written as $\mathds{1}(\mathbf{x}^{V}\in S)$. Thus, we have that
@@
$=\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{V^{\perp}}^{S}}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq y]\;,$
where $D_{V^{\perp}}^{S}$ denotes the marginal distribution on $V^{\perp}$ obtained by averaging $D$ over $V$ while conditioning on the event $\mathbf{x}^{V}\in S$. Similarly, we have $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{V^{\perp}}^{S}}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq i]>C\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{V^{\perp}}^{S}}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq y]+\epsilon$ for all $i\in[K]$. Moreover, by the properties of the standard Gaussian distribution, the $\mathbf{x}$-marginal of $D_{V^{\perp}}^{S}$ is a standard Gaussian.
As a result, by applying Proposition 2.4 for $f_{\mathcal{G}_{S}}$, we have that there exist $i,j\in[K],i\neq j$ such that $(\widehat{\mathbf{w}}^{(i,j)})^{\perp V}\cdot\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D^{S}_{V^{\perp}}}[\mathbf{x}\mathds{1}(y=i)]\geq\mathrm{poly}(1/K,\epsilon)$. Furthermore, note that by averaging over $V$ conditioning on $S$ we have that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{x}^{\perp V}\mathds{1}(y=i)\mid\mathbf{x}\in S]=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}^{S}}[\mathbf{x}\mathds{1}(y=i)]$. Therefore, there exist $i,j\in[K],i\neq j$, such that $(\widehat{\mathbf{w}}^{(i,j)})^{\perp V}\cdot\mathbf{u}^{(S,l)}\geq\mathrm{poly}(1/K,\epsilon)$ for some $l\in[K]$.
@@
Let $p(\epsilon,1/K)$ be a sufficiently small polynomial in $\epsilon$ and $1/K$. If $N=d(k/\epsilon^{\prime})^{Ck}(K/\epsilon)^{C}\log(1/\delta)$ for a sufficiently large universal constant $C>0$, then with probability $1-\delta$ we have that $\left\|(\widehat{\mathbf{u}}^{(S,i)}-{\mathbf{u}}^{(S,i)})^{\perp V}\right\|\leq p(\epsilon,1/K)$ and $\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]\geq\operatorname*{\mathbf{Pr}}_{D}[S]/2$ for all $S\in\mathcal{S},i\in[K]$.
@@
First note that by the definition of an approximating partition (see Definition 2.2) and A.1, for all $S\in\mathcal{S}$ we have that $\operatorname*{\mathbf{Pr}}_{D}[S]=(\epsilon^{\prime}/k)^{\Omega(k)}$ and $\lvert\mathcal{S}\rvert=(k/\epsilon^{\prime})^{O(k)}$. Therefore, by the union bound and Hoeffding’s inequality, if $N\geq(k/\epsilon^{\prime})^{Ck}\log(\lvert\mathcal{S}\rvert/\delta)=(k/\epsilon^{\prime})^{O(k)}\log(1/\delta)$ for a sufficiently large absolute constant $C>0$, with probability at least $1-\delta$ we have that $\lvert\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]-\operatorname*{\mathbf{Pr}}_{D}[S]\rvert\leq\operatorname*{\mathbf{Pr}}_{D}[S]/2$ for all $S\in\mathcal{S}$. Hence, $\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]\geq\operatorname*{\mathbf{Pr}}_{D}[S]/2$, i.e., the number of samples that fall in each set $S\in\mathcal{S}$ is at least $N\operatorname*{\mathbf{Pr}}_{D}[S]/2=N(\epsilon^{\prime}/k)^{\Omega(k)}$.
Furthermore, note that by the definitions of ${\mathbf{u}}^{(S,i)}$ and $\widehat{\mathbf{u}}^{(S,i)}$ it holds that
$(\widehat{\mathbf{u}}^{(S,i)}-{\mathbf{u}}^{(S,i)})^{\perp V}$ $=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim\widehat{D}}[\mathbf{x}^{\perp V}\mathds{1}(y=i)\mid\mathbf{x}\in S]-\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{x}^{\perp V}\mathds{1}(y=i)\mid\mathbf{x}\in S]$
Therefore, to prove our concentration result, it suffices to show that when conditioning on $\mathbf{x}\in S$ the random vector $\mathbf{x}^{\perp V}\mathds{1}(y=i)$ with $(\mathbf{x},y)\sim D$ is subgaussian in every unit direction with subgaussian norm bounded by a constant. To see this, note that whether $\mathbf{x}$ belongs in $S$ depends only on the projection of $\mathbf{x}$ onto $V$, and $\mathbf{x}^{\perp V}$ is independent of $\mathbf{x}^{V}$. Now fix a unit vector $\mathbf{v}\in\mathbb{R}^{d}$, we have that for any $p\geq 1$
$\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}\left[\left(\mathbf{v}\cdot\mathbf{x}^{\perp V}\mathds{1}(y=i)-\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{v}\cdot\mathbf{x}^{\perp V}\mathds{1}(y=i)\mid\mathbf{x}\in S]\right)^{p}\mid S\right]^{1/p}$ $\lesssim\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}\left[\left(\mathbf{v}\cdot\mathbf{x}^{\perp V}\mathds{1}(y=i)\right)^{p}\mid S\right]^{1/p}$
$\leq\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}\left[\left(\mathbf{v}\cdot\mathbf{x}^{\perp V}\right)^{p}\right]^{1/p}$
$\leq\operatorname*{\mathbf{E}}_{z\sim\mathcal{N}(0,1)}\left[(z)^{p}\right]^{1/p}\lesssim\sqrt{p}\;,$
@@
where in the first inequality we used the triangle and Jensen’s inequalities, while in the second we exploited the independence of $\mathbf{x}^{V}$ and $\mathbf{x}^{\perp V}$ and the fact that the indicator function is bounded by 1. Therefore $\mathbf{x}^{\perp V}\mathds{1}(y=i)$ is subgaussian. Hence, by the use of a standard covering argument it follows that if the number of samples that falls in each discretization set $S$ is at least $d\mathrm{poly}(K,1/\epsilon)\log(K/\delta)$ for a sufficiently large polynomial, then $\left\|(\widehat{\mathbf{u}}^{(S,i)}-{\mathbf{u}}^{(S,i)})^{\perp V}\right\|\leq p(\epsilon,1/K)$ for all $i\in[K]$.
@@
Combining the above results along with the union bound gives us that there exists an $N=d(k/\epsilon^{\prime})^{O(k)}(K/\epsilon)^{O(1)}\log(1/\delta)$ that suffices. ∎
@@
Applying Claims 2.11 and 2.12, we have that if $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{\mathcal{S}}(\mathbf{x})\neq y]>(C+1)\mathrm{OPT}+\epsilon$ then there exists a subset $\mathcal{T}\subseteq\mathcal{S}$ of the $\epsilon^{\prime}$-approximating partition $\mathcal{S}$ with $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}_{D}[S]=\Omega(\epsilon)$ such that for all $S\in\mathcal{T}$ we have $(\widehat{\mathbf{w}}^{(i,j)})^{\perp V}\cdot\mathbf{u}^{(S,l)}\geq\mathrm{poly}(1/K,\epsilon)$ for some $i,j,l\in[K],i\neq j$. Moreover, if $N=d(k/\epsilon^{\prime})^{C^{\prime}k}(K/\epsilon)^{C^{\prime}}\log(1/\delta)$ for a sufficiently large universal constant $C^{\prime}>0$, using the concentration result of Claim 2.13, we have that the empirical moments of each set satisfy the same guarantees, i.e., $(\widehat{\mathbf{w}}^{(i,j)})^{\perp V}\cdot\widehat{\mathbf{u}}^{(S,l)}\geq\mathrm{poly}(1/K,\epsilon)$, for some $i,j,l\in[K],i\neq j,$ for all $S\in\mathcal{T}$. Furthermore, note that by Claim 2.13 we also have that $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]=\Omega(\epsilon)$.
@@
Therefore, there exists a subset $\mathcal{T}^{\prime}\subseteq\mathcal{T}$ and $i,j\in[K]$ with $\sum_{S\in\mathcal{T}^{\prime}}\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]\geq\Omega(\epsilon/K^{2})$ such that for all $S\in\mathcal{T}^{\prime}$ we have that $(\widehat{\mathbf{w}}^{(i,j)})^{\perp V}\cdot\widehat{\mathbf{u}}^{(S,l)}\geq\mathrm{poly}(1/K,\epsilon)$ for some $l\in[K]$. Hence, for some $i\neq j\in[K]$ we have that
@@
Since the set $S$ depends only on the projection $\mathbf{x}^{V}$, we have $(\mathbf{u}^{(S,i)})^{\perp V}=\operatorname*{\mathbf{E}}_{\mathbf{x}^{\perp V}}[\mathbf{x}^{\perp V}\operatorname*{\mathbf{E}}_{\mathbf{x}^{V}}[\mathds{1}(y=i)\mid\mathbf{x}\in S]]$. Therefore, we can bound the norm of $(\mathbf{u}^{(S,i)})^{\perp V}$ as follows:
@@
where we used the Cauchy-Schwarz inequality and the dual characterization of the $\ell_{2}$ norm. In the last inequality we used that $\mathbf{x}^{\perp V}$ is a standard Gaussian, and henceforth the first expectation is at most $1$, and the second expectation is at most $1$ as well because $\mathds{1}(y=i)$ is bounded by $1$. By the triangle inequality, it follows that
@@
Hence, by Claim B.1, there exists a unit eigenvector $\mathbf{v}$ of $\widehat{\mathbf{U}}$ with eigenvalue at least $\mathrm{poly}(\epsilon/K)$ (for a sufficiently small polynomial), such that $(\widehat{\mathbf{w}}^{(i,j)})^{\perp V}\cdot\mathbf{v}\geq\mathrm{poly}(1/K,\epsilon)$, for some $i\neq j\in[K]$. Moreover, the number of such eigenvectors is at most $2K/\mathrm{poly}(1/K,\epsilon)=\mathrm{poly}(K,1/\epsilon)$. This completes the proof of Proposition 2.8. ∎
Before proving our main result of this section, Theorem 2.1, we first show that if we have learned one direction of our target model to sufficiently high accuracy, then we may, without loss of generality, restrict the model to the remaining directions. Specifically, we establish the following: